a) The speed of sound through steel is 5,200 m/s. This is 2,520 m/s faster than the speed of sound through silver. Write and solve an equation to find the speed of sound through silver.

b) Find the speed of sound in a steel rod is Young’s modulus of steel is 2.07x1011Pa and its density is 78000kg/m3

c) What is the speed of sound in air when the temperature is 20°C? y=1.4, R=8.314J/mol K and the molar mass of air is 29 g/mol


1
Expert's answer
2021-09-23T14:41:24-0400

(a) "c_{silver}=5200-2520=2680\ (m/s)"


(b) "c_{steel}=\sqrt{E/\rho_{steel}}=\sqrt{207\cdot10^9/7800}=5152\ (m/s)"


(c) "c=\sqrt{\gamma RT/M}=\sqrt{1.4\cdot8.314\cdot293/0.029}=343\ (m/s)"